# SIC-POVMs and orders of real quadratic fields

### v1

@@ 3. Fields generated by SICs
In this section, we define several number fields attached to arbitrary SICs. Many of these fields have been studied before, but they are often discussed within the context of assuming various conjectures about SICs. Here, we focus on relationships between “SIC fields” that can be proven rigorously, either in general (Section 3.1) or in the case of Weyl–Heisenberg SICs (Section 3.2).
@@ 3.1. Fields generated by general SICs
The fields we will study here are each associated to an individual SIC $S$ and depend only on data computed from the orthogonal projections $\{\Pi_{i}:1\leq i\leq d^{2}\}$ defining the SIC.
@@
Let ${\mathbf{\Pi}}(S)=\{\frac{1}{d}\Pi_{1},\ldots,\frac{1}{d}\Pi_{d^{2}}\}$ be an arbitrary SIC in dimension $d$, with each $\Pi_{i}=\mathbf{v}_{i}\mathbf{v}_{i}^{\dagger}$ for a unit vector $\mathbf{v}_{i}$. (Thus each $\mathbf{v}_{i}$ is a unit fiducial vector, unique up to multiplication by a unit scalar $e^{i\theta}$.)
@@
    The ratio SIC-field $F^{\rm vec}\!\left(S\right)$ is the field extension of $\mathbb{Q}$ generated by all the ratios $v_{ij}/v_{ik}$ (for $v_{ik}\neq 0$) of the entries of the $d^{2}$ fiducial vectors $\mathbf{v}_{i}$ of $S$.
@@
    The triple product SIC-field $F^{\rm trip}\!\left(S\right)$ is the field extension of $\mathbb{Q}$ generated by all the “triple products” $\operatorname{Tr}(\Pi_{i}\Pi_{j}\Pi_{k})$.
@@
    The unitary invariant SIC-field $F^{\rm inv}\!\left(S\right)$ is the field extension of $\mathbb{Q}$ generated by all the unitary invariants $\operatorname{Tr}(\Pi_{i_{1}}\Pi_{i_{2}}\cdots\Pi_{i_{n}})$ for $n\in\mathbb{N}$.
@@
    The projection SIC-field $F^{\rm proj}\!\left(S\right)$ is the field extension of $\mathbb{Q}$ generated by all the entries of the $d^{2}$ Hermitian projection matrices $\Pi_{j}$.
@@
    The extended projection SIC-field $F^{\rm eproj}\!\left(S\right):=F^{\rm proj}\!\left(S\right)\!(\xi_{d})$, adjoining $\xi_{d}=-e^{\frac{\pi i}{d}}$.
A priori, these fields could be transcendental extensions of $\mathbb{Q}$, and this does occur in dimension $d=3$.
@@
These fields have previously been studied with various names and notations, which we now review.
@@
    The field $F^{\rm vec}\!\left(S\right)$ is the SIC field $\mathbb{Q}[S]$ attached to a general SIC $S$ by Appleby, Flammia, McConnell, and Yard [10, p. 212 bottom], as used in Thm. 1 and Conj. 2 of that paper. By definition, this field is independent of scaling factors in the $\mathbf{v}_{i}$. It is determined by ${\mathbf{\Pi}}(S)$ because the ratios
@@
    Prop. 3 of [10] states that, for a Weyl–Heisenberg SIC, the ratio SIC-field is the same for any SIC $S^{\prime}\in[S]$. The presented proof outline [10, Sec. 7.1] appears to assume some conjectural equivalences and characterizations of other related fields.
@@
    Concerning the triple product SIC-field $F^{\rm trip}\!\left(S\right)$, the fact that a SIC is determined up to unitary equivalence by its triple products was proved in 2011 by Appleby, Flammia, and Fuchs [3, Thm. 3]. We show unconditionally that $F^{\rm trip}\!\left(S\right)$ is an invariant of the geometric $\operatorname{PEC}(d)$-orbit of a Weyl–Heisenberg SIC $S$; see 3.4.
@@
    The unitary invariant SIC-field $F^{\rm inv}\!\left(S\right)$ is generated by the full set of projective unitary invariants of a general finite set of lines in $\mathbb{C}^{d}$; see [15] and [46, Ch. 8, Thm. 8.2]. We show in 3.3 that $F^{\rm trip}\!\left(S\right)=F^{\rm inv}\!\left(S\right)$.
@@
    The projection SIC-field $F^{\rm proj}\!\left(S\right)$, in the special case that $S$ is a Weyl–Heisenberg SIC, coincides with the field $L$ appearing in [10, Sec. 7], defined as the smallest subfield of $\mathbb{C}$ generated by the entries $v_{j}\overline{v}_{k}$ of unit vectors generating $\Pi_{\mathbf{v}}$ taken over all $d^{2}$ projectors $\Pi_{\mathbf{v}}$ of a Weyl–Heisenberg SIC.
@@
    The extended projection SIC-field $F^{\rm eproj}\!\left(S\right)$ is the SIC field $\mathbb{E}$ as defined by Appleby, Flammia, McConnell, and Yard [9, Sec. 4] for any SIC projector $S$. This definition of the SIC field is also used in [7]. Appleby, Flammia, McConnell, and Yard [9] state (without proof) that the field $\mathbb{E}$ is well-defined on the $\operatorname{PEC}(d)$-orbit $[S]$ of a Weyl–Heisenberg SIC $S$. We prove so unconditionally in 3.8.
@@
Proof of (1): Write $\Pi_{i}=\mathbf{v}_{i}\mathbf{v}_{i}^{\dagger}$, where $\mathbf{v}_{i}$ has entries $v_{ij}$ and $\left|\mathbf{v}_{i}\right|^{2}=1$. The nonzero entries of $\Pi_{i}$ are of the form $v_{ij}\overline{v}_{ik}$ for $v_{ij}\neq 0$, $v_{ik}\neq 0$ These entries can be expressed as
$v_{ij}\overline{v}_{ik}=\frac{v_{ij}\overline{v}_{ik}}{\left|\mathbf{v}_{i}\right|^{2}}=\frac{1}{\sum_{\ell=1}^{d}(v_{i\ell}/v_{ij})\overline{(v_{i\ell}/v_{ik})}}\in F^{\rm vec}\!\left(S\right)\overline{F^{\rm vec}\!\left(S\right)}.$ (3.2)
Thus, $F^{\rm proj}\!\left(S\right)\subseteq F^{\rm vec}\!\left(S\right)\overline{F^{\rm vec}\!\left(S\right)}$. To show the reverse inclusion, note that for $v_{ik}\neq 0$,
$v_{ij}/v_{ik}=\frac{v_{ij}\overline{v}_{ik}}{v_{ik}\overline{v}_{ik}}\mbox{ and }\overline{(v_{ij}/v_{ik})}=\frac{v_{ik}\overline{v}_{ij}}{v_{ik}\overline{v}_{ik}}.$ (3.3)
Proof of (2): First we show the equality $F^{\rm trip}\!\left(S\right)=F^{\rm inv}\!\left(S\right)$. Clearly $F^{\rm trip}\!\left(S\right)\subseteq F^{\rm inv}\!\left(S\right)$. To show the reverse inclusion, note that the $\Pi_{i}$ form a basis for the space of $d\times d$ complex matrices, and express
@@
for structure constants $\alpha_{ij}^{\ell}\in\mathbb{C}$. We have
@@
Thus, $\alpha_{ij}^{k}=\frac{d+1}{d}\operatorname{Tr}(\Pi_{i}\Pi_{j}\Pi_{k})-\frac{d(\delta_{ij}d+1)}{d+1}$, and the field generated over $\mathbb{Q}$ by the structure constants is $F^{\rm trip}\!\left(S\right)$. Moreover, using (3.4) repeatedly, we can express any product $\Pi_{i_{1}}\Pi_{i_{2}}\cdots\Pi_{i_{n}}$ in the $\left\{\Pi_{\ell}\right\}_{\ell}$-basis with coefficients that are integer polynomials in the structure constants. Thus, since $\operatorname{Tr}(\Pi_{\ell})=1$, $\operatorname{Tr}(\Pi_{i_{1}}\Pi_{i_{2}}\cdots\Pi_{i_{n}})$ is also an integer polynomials in the structure constants and is thus contained in $F^{\rm trip}\!\left(S\right)$.
Finally, the inclusion $F^{\rm inv}\!\left(S\right)\subseteq F^{\rm proj}\!\left(S\right)$ of (2) follows because $\operatorname{Tr}(\Pi_{i_{1}}\Pi_{i_{2}}\cdots\Pi_{i_{n}})$ is an integer polynomial in the entries of the various $\Pi_{j}$. ∎
@@
Let $S=\{\mathbb{C}{\mathbf{v}}_{i}:1\leq i\leq d^{2}\}$ be an arbitrary line-SIC, and set ${\mathbf{\Pi}}(S)=\{\frac{1}{d}\Pi_{1},\ldots,\frac{1}{d}\Pi_{d^{2}}\}$, with $\Pi_{i}={\mathbf{v}}_{i}{\mathbf{v}}_{i}^{\dagger}$. Then ${\mathbf{\Pi}}(US)=\{\frac{1}{d}\Pi_{1}^{\prime},\ldots,\frac{1}{d}\Pi_{d^{2}}^{\prime}\}$. Write $U=CV$, where $C$ is either the identity or complex conjugation, and $V\in\mathbf{U}(d)$; then either $\Pi_{i}^{\prime}=V^{-1}\Pi_{i}V$, or $\Pi_{i}^{\prime}=V^{-1}\Pi_{i}^{\top}V$. In the first case, $\operatorname{Tr}(\Pi_{i}^{\prime}\Pi_{j}^{\prime}\Pi_{k}^{\prime})=\operatorname{Tr}(V^{-1}\Pi_{i}\Pi_{j}\Pi_{k}V)=\operatorname{Tr}(\Pi_{i}\Pi_{j}\Pi_{k})$. In the second case, $\operatorname{Tr}(\Pi_{i}^{\prime}\Pi_{j}^{\prime}\Pi_{k}^{\prime})=\operatorname{Tr}\!\left(V^{-1}\Pi_{i}^{\top}\Pi_{j}^{\top}\Pi_{k}^{\top}V\right)=\operatorname{Tr}\!\left((\Pi_{k}\Pi_{j}\Pi_{i})^{\top}\right)=\operatorname{Tr}(\Pi_{k}\Pi_{j}\Pi_{i})$. So $F^{\rm trip}\!\left(US\right)=F^{\rm trip}\!\left(S\right)$ in both cases. ∎
@@
Proof of (1): The implication $F^{\rm vec}\!\left(S\right)=F^{\rm proj}\!\left(S\right)$ follows from 3.3(1).
Proof of (2): If $K=F^{\rm proj}\!\left(S\right)$ is Galois over $\mathbb{Q}$, then $F^{\rm eproj}\!\left(S\right)$ is the compositum $KL$, with $L=\mathbb{Q}(e^{\frac{\pi i}{d}})$ an abelian extension of $\mathbb{Q}$, so the compositum is a Galois extension of $\mathbb{Q}$. ∎
@@ 3.2. Fields generated by Weyl–Heisenberg SICs
We now derive some properties of the field $F^{\rm vec}\!\left(S\right)$, $F^{\rm trip}\!\left(S\right)$ and $F^{\rm proj}\!\left(S\right)$ in the special case of Weyl–Heisenberg SICs.
Lemma 3.6.
Let $S$ be a Weyl–Heisenberg line-SIC $S=S([{\mathbf{v}}])$ of dimension $d$. Then:
@@
    $F^{\rm proj}\!\left(S\right)$ contains a $d$-th root of unity $\zeta_{d}=e^{\frac{2\pi i}{d}}$.

### v3

@@ 3. Fields generated by SICs
In this section, we define several number fields attached to arbitrary SICs. Many of these fields have been studied before, often within the context of assuming various conjectures about SICs. Here, we focus on relationships between “SIC fields” that can be proven rigorously, either in general (Section 3.1) or in the case of Weyl–Heisenberg SICs (Section 3.2).
@@ 3.1. Fields generated by general SICs
The fields we study here are each associated to an individual SIC $S$ and depend only on data computed from the orthogonal projections $\{\Pi_{i}:1\leq i\leq d^{2}\}$ defining the SIC.
@@
Let ${\mathbf{\Pi}}(S)=\{\frac{1}{d}\Pi_{1},\ldots,\frac{1}{d}\Pi_{d^{2}}\}$ be an arbitrary SIC in dimension $d$, with each $\Pi_{i}=\mathbf{v}_{i}\mathbf{v}_{i}^{\dagger}$ for a unit vector $\mathbf{v}_{i}$. (Thus each $\mathbf{v}_{i}$ is unique up to multiplication by a unit scalar.)
@@
    The ratio SIC field $F^{\rm vec}\!\left(S\right)$ is the field extension of $\mathbb{Q}$ generated by all the ratios $v_{ij}/v_{ik}$ (for $v_{ik}\neq 0$) of the entries of the $d^{2}$ fiducial vectors $\mathbf{v}_{i}$ of $S$.
@@
    The triple product SIC field $F^{\rm trip}\!\left(S\right)$ is the field extension of $\mathbb{Q}$ generated by all the triple products $\operatorname{Tr}(\Pi_{i}\Pi_{j}\Pi_{k})$ (also called $3$-vertex Bargmann invariants [46, Sec. 8.2]).
@@
    The unitary invariant SIC field $F^{\rm inv}\!\left(S\right)$ is the field extension of $\mathbb{Q}$ generated by all the unitary invariants $\operatorname{Tr}(\Pi_{i_{1}}\Pi_{i_{2}}\cdots\Pi_{i_{n}})$ for $n\in\mathbb{N}$ (also called $n$-vertex Bargmann invariants [46, Sec. 8.2]).
@@
    The projection SIC field $F^{\rm proj}\!\left(S\right)$ is the field extension of $\mathbb{Q}$ generated by all the entries of the $d^{2}$ Hermitian projection matrices $\Pi_{j}$.
@@
    The extended projection SIC field $F^{\rm eproj}\!\left(S\right):=F^{\rm proj}\!\left(S\right)\!(\xi_{d})$, adjoining $\xi_{d}=-e^{\frac{\pi i}{d}}$.
A priori, these fields could be transcendental extensions of $\mathbb{Q}$, as happens in dimension $d=3$.
@@
These fields have previously been studied with various names and notations.
@@
    The field $F^{\rm vec}\!\left(S\right)$ is the SIC field $\mathbb{Q}[S]$ attached to a general SIC $S$ by Appleby, Flammia, McConnell, and Yard [9, p. 212 bottom], as used in Thm. 1 and Conj. 2 of that paper. By definition, this field is independent of scaling factors in the $\mathbf{v}_{i}$. It is determined by ${\mathbf{\Pi}}(S)$, because the ratios
@@
    It is stated as [9, Prop. 3] that, for a Weyl–Heisenberg SIC, the ratio SIC field is the same for any SIC $S^{\prime}\in[S]$. The presented proof outline [9, Sec. 7.1] appears to assume some conjectural equivalences and characterizations of other related fields.
@@
    Concerning the triple product SIC field $F^{\rm trip}\!\left(S\right)$, the fact that a SIC is determined up to unitary equivalence by its triple products was proved in 2011 by Appleby, Flammia, and Fuchs [3, Thm. 3]; see also [46, Cor. 8.1]. We show unconditionally that $F^{\rm trip}\!\left(S\right)$ is an invariant of the geometric $\operatorname{PEC}(d)$-orbit of a Weyl–Heisenberg SIC $S$; see 3.4.
@@
    The unitary invariant SIC field $F^{\rm inv}\!\left(S\right)$ is generated by the full set of projective unitary invariants of a general finite set of lines in $\mathbb{C}^{d}$; see [14] and [46, Ch. 8, Thm. 8.2]. We show in 3.3 that $F^{\rm trip}\!\left(S\right)=F^{\rm inv}\!\left(S\right)$.
@@
    The projection SIC field $F^{\rm proj}\!\left(S\right)$, in the special case that $S$ is a Weyl–Heisenberg SIC, coincides with the field $L$ appearing in [9, Sec. 7], defined as the smallest subfield of $\mathbb{C}$ generated by the entries $v_{j}\overline{v}_{k}$ of unit vectors generating $\Pi_{\mathbf{v}}$ taken over all $d^{2}$ projectors $\Pi_{\mathbf{v}}$ of a Weyl–Heisenberg SIC.
@@
    The extended projection SIC field $F^{\rm eproj}\!\left(S\right)$ is the SIC field $\mathbb{E}$ as defined by Appleby, Flammia, McConnell, and Yard [8, Sec. 4] for any SIC projector $S$. This definition of the SIC field is also used in [6] and is treated in detail in [46, Sec. 14.20]. Appleby, Flammia, McConnell, and Yard [8] state that the field $\mathbb{E}$ is well-defined on the $\operatorname{PEC}(d)$-orbit $[S]$ of a Weyl–Heisenberg SIC $S$. We supply a proof in 3.8.
@@
Proof of (1). Write $\Pi_{i}=\mathbf{v}_{i}\mathbf{v}_{i}^{\dagger}$, where $\mathbf{v}_{i}$ has entries $v_{im}$ and $\left|\mathbf{v}_{i}\right|^{2}=1$, for $1\leq i\leq d^{2}$. The entries of $\Pi_{i}$ are of the form $v_{im}\overline{v}_{in}$ for $10\leq m,n\leq d-$ 1 When both $v_{im}\neq 0$, $v_{in}\neq 0$, these entries can be expressed as
$v_{im}\overline{v}_{in}=\frac{v_{im}\overline{v}_{in}}{\left|\mathbf{v}_{i}\right|^{2}}=\frac{1}{\sum_{p=0}^{d-1}(v_{ip}/v_{im})\overline{(v_{ip}/v_{in})}}\in F^{\rm vec}\!\left(S\right)\overline{F^{\rm vec}\!\left(S\right)},$ (3.2)
where $10\leq m,n,p\leq d-$ 1. Thus, $F^{\rm proj}\!\left(S\right)\subseteq F^{\rm vec}\!\left(S\right)\overline{F^{\rm vec}\!\left(S\right)}$. To show the reverse inclusion, note that for $v_{ik}\neq 0$,
$v_{im}/v_{in}=\frac{v_{im}\overline{v}_{in}}{v_{im}\overline{v}_{im}}\mbox{ and }\overline{(v_{im}/v_{in})}=\frac{v_{in}\overline{v}_{im}}{v_{in}\overline{v}_{in}}.$ (3.3)
Proof of (2). First, the inclusion $F^{\rm inv}\!\left(S\right)\subseteq F^{\rm proj}\!\left(S\right)$ of (2) follows because $\operatorname{Tr}(\Pi_{i_{1}}\Pi_{i_{2}}\cdots\Pi_{i_{n}})$ is an integer polynomial in the entries of the various $\Pi_{j}$.
Second, we show the equality $F^{\rm trip}\!\left(S\right)=F^{\rm inv}\!\left(S\right)$. Clearly $F^{\rm trip}\!\left(S\right)\subseteq F^{\rm inv}\!\left(S\right)$. To show the reverse inclusion, note that the $\Pi_{i}$ form a basis for the space of $d\times d$ complex matrices, and express
@@
for structure constants $\alpha_{ij}^{\ell}\in\mathbb{C}$, having $1\leq i,j,\ell\leq d^{2}$. We have
@@
Thus, $\alpha_{ij}^{k}=\frac{d+1}{d}\operatorname{Tr}(\Pi_{i}\Pi_{j}\Pi_{k})-\frac{d(\delta_{ij}d+1)}{d+1}$, and the field generated over $\mathbb{Q}$ by the structure constants is $F^{\rm trip}\!\left(S\right)$. Moreover, using (3.4) repeatedly, we can express any product $\Pi_{i_{1}}\Pi_{i_{2}}\cdots\Pi_{i_{n}}$ in the $\left\{\Pi_{\ell}\right\}_{\ell}$-basis with coefficients that are integer polynomials in the structure constants. Thus, since $\operatorname{Tr}(\Pi_{\ell})=1$, $\operatorname{Tr}(\Pi_{i_{1}}\Pi_{i_{2}}\cdots\Pi_{i_{n}})$ is also an integer polynomial in the structure constants and is thus contained in $F^{\rm trip}\!\left(S\right)$. ∎
@@
Write $S=\{\mathbb{C}{\mathbf{v}}_{i}:1\leq i\leq d^{2}\}$, and set ${\mathbf{\Pi}}(S)=\{\frac{1}{d}\Pi_{1},\ldots,\frac{1}{d}\Pi_{d^{2}}\}$ with $\Pi_{i}={\mathbf{v}}_{i}{\mathbf{v}}_{i}^{\dagger}$. Then set ${\mathbf{\Pi}}(US)=\{\frac{1}{d}\Pi_{1}^{\prime},\ldots,\frac{1}{d}\Pi_{d^{2}}^{\prime}\}$. Write $U=CV$, where $C$ is identity or complex conjugation, and $V\in\mathbf{U}(d)$; then either $\Pi_{i}^{\prime}=V^{-1}\Pi_{i}V$, or $\Pi_{i}^{\prime}=V^{-1}\Pi_{i}^{\top}V$. In the first case, $\operatorname{Tr}(\Pi_{i}^{\prime}\Pi_{j}^{\prime}\Pi_{k}^{\prime})=\operatorname{Tr}(V^{-1}\Pi_{i}\Pi_{j}\Pi_{k}V)=\operatorname{Tr}(\Pi_{i}\Pi_{j}\Pi_{k})$. In the second case,
$\operatorname{Tr}(\Pi_{i}^{\prime}\Pi_{j}^{\prime}\Pi_{k}^{\prime})=\operatorname{Tr}\!\left(V^{-1}\Pi_{i}^{\top}\Pi_{j}^{\top}\Pi_{k}^{\top}V\right)=\operatorname{Tr}\!\left((\Pi_{k}\Pi_{j}\Pi_{i})^{\top}\right)=\operatorname{Tr}(\Pi_{k}\Pi_{j}\Pi_{i}).$ (3.9)
So $F^{\rm trip}\!\left(US\right)=F^{\rm trip}\!\left(S\right)$ in both cases. ∎
@@
The implication $F^{\rm vec}\!\left(S\right)=F^{\rm proj}\!\left(S\right)$ follows from 3.3(1). Moreover, if $K=F^{\rm proj}\!\left(S\right)$ is Galois over $\mathbb{Q}$, then $F^{\rm eproj}\!\left(S\right)$ is the compositum $KL$, with $L=\mathbb{Q}(e^{\frac{\pi i}{d}})$ an abelian extension of $\mathbb{Q}$, so the compositum is a Galois extension of $\mathbb{Q}$. ∎
@@ 3.2. Fields generated by Weyl–Heisenberg SICs
We now derive some properties of the fields $F^{\rm vec}\!\left(S\right)$, $F^{\rm trip}\!\left(S\right)$, $F^{\rm proj}\!\left(S\right)$, and $F^{\rm eproj}\!\left((\right)S)$ in the special case of Weyl–Heisenberg SICs.
Proposition 3.6.
Let $S$ be a Weyl–Heisenberg line-SIC $S=S([{\mathbf{v}}])$ of dimension $d\geq 2$. Then:
@@
    The field $F^{\rm trip}\!\left(S\right)$ contains the $d$-th root of unity $\zeta_{d}=e^{\frac{2\pi i}{d}}$.
  • (2)
    The field $F^{\rm proj}\!\left(S\right)\subseteq L(S)=F^{\rm trip}\!\left(S\right)\!({\sqrt[d]{\beta_{1}},\sqrt[d]{\beta_{2}}})$, for some $\beta_{1},\beta_{2}\in F^{\rm trip}\!\left(S\right)$. $L(S)$ is an abelian extension of $F^{\rm trip}\!\left(S\right)$ whose degree divides $d^{2}$, hence $[F^{\rm proj}\!\left(S\right):F^{\rm trip}\!\left(S\right)]$ divides $d^{2}$.